Give the exact and approximate solutions to three decimal places.
y^2-10y+25=100
equation is ax^2+bx+c I believe ??? I am confused because there are y's and 10 is negative...

Answer by stanbon(75887)

Give the exact and approximate solutions to three decimal places.
y^2-10y+25=100
---
Factor:
(y-5)^2 - 100 = 0
---
Factor
[(y-5)-10][(y-5)+10] = 0
---
y-5 = 10 or y-5 = -10
y = 15 or y = -5
=======================
Cheers,
Stan H.
==================